an 1000kg car traveling at 80 m/s takes 19m to stop under full breaking. the same car under similar road conditions, traveling at 160m/s, takes ___________ m to stop under full breaking.
Answer #1

here,

initial speed , u1 = 80 m/s

distance taken to stop , d1 = 19 m

let the accelration be a

using third equation of motion

v1^2 - u1^2 = 2 * d1 * a

0 - 80^2 = 2 * 19 * a

a = - 168.42 m/s^2

when

initial speed , u2 = 160 m/s

let the distance taken to stop be d2

using third equation of motion

v2^2 - u2^2 = 2 * d2 * a

0 - 160^2 = - 2 * d2 * 168.42

d2 = 76 m

the distance taken to stop is 76 m